# Quantum Algorithms in a Superposition of Spacetimes

### v1

@@ Abstract
We show that a quantum computer with the ability to create a superposition of unitary orders is able to solve in polynomial time two of the fundamental problems in computer science: The Graph Isomorphism Problem ($\mathsf{GI}$) and the Gap Closest Vector Problem ($\mathsf{GapCVP}$), with gap $O\left(n^{2}\right)$. These problems are believed by experts to be hard to solve for a regular quantum computer. Interestingly, our model does not seem overpowered, and we found no obvious way to solve entire complexity classes that are considered hard in computer science, like the classes NP and SZK.
@@ 1 Introduction
Formulating an experimentally-verified theory of physics that generalizes both QM and GR, is one of the grand challenges of physics. Such a theory, usually referred to as a theory of Quantum Gravity, will model the interaction between quantum effects and gravitational force; an interaction which is currently ignored in most analyses. The two most popular candidates as theories of quantum gravity are String Theory and Loop Quantum Gravity, however, numerous other proposals exist [${}^{+}$22].
@@
The goal of this paper is to suggest and define a computational model for a quantum computer with the added ability of generating a superposition of spacetime geometries, and to analyze its power. The rest of the Introduction is as follows. In Section 1.1 we provide a brief overview of the physics literature on a superposition of spacetime geometries, and then explain how it serves as inspiration for our computational model. This part deals with physics, it is self contained and can be skipped for the readers that are interested only in the complexity theoretical aspects of this work. In Section 1.2 we present our computational model and its connection to both standard quantum computation and previous models in the quantum information theory. In Section 1.3 we present our complexity theoretical results for the model.
@@ 1.1 Gravity, Superposition and Entanglement
Before we explain what a quantum spacetime may be able to do, we refer to one ability of a classical spacetime. GR postulates the effects of gravitational time dilation: The structure of spacetime determines the pace at which time flows. More precisely, gravitational time dilation is the effect of time moving slower in a region of space with greater gravitational potential: For some location $s\in\mathbb{R}^{3}$ in space, the time in $s$ slows down proportionally to how close $s$ is to a massive object, and the mass of that object. According to this logic, a quantum spacetime may enable a superposition of spacetime curvatures, which in turn induces a superposition of time dilation effects, on the same set of events.
@@
  • The last detail is the only one that uses effects from GR, specifically, gravitational time dilation. The system $S_{M}$ is located with respect to the second system $S_{U}$ such that, even though the systems are spatially disjoint in coordinates, if $M$ is at location $L_{0}$, which is closer to clock $C_{1}$, then $C_{1}$ ticks slower and $C_{0}$ completes first, and if $M$ is at location $L_{1}$, which is closer to $C_{0}$, then $C_{0}$ ticks slower and $C_{1}$ completes first.
Imagine we (1) execute $S_{M}$ and generate a uniform superposition of the two possible locations for $M$, followed by (2) executing $S_{U}$, the clocked execution of the unitaries $U_{0}$, $U_{1}$. We can analyze what happens in each of the branches of the superposition, separately, assuming as if things are classical and there is no superposition. If $M$ is at $L_{0}$, the state in $R$ at the end of the execution is $U_{1}U_{0}|{\psi}\rangle$, and if $M$ is at $L_{1}$ then the state in $R$ at the end of execution is $U_{0}U_{1}|{\psi}\rangle$. The question becomes one in quantum gravity, when we ask what happens not in each branch of the superposition separately, but as a full joint system, in superposition. Experimental physics does not know the answer to this question.
The above question may translate to more basic questions about gravity, superposition and entanglement. Specifically, the first question is whether a massive object can be maintained in a quantum superposition, and the second question is whether entanglement can be created through gravity alone. More elaborately, on one side, the gravitational decoherence hypothesis [Dió89, Pen96, Dio14, Dio87] says that a quantum superposition of a massive object will fundamentally decohere, turning into a probabilistic mixture of classical states. On the other end, gravitationally-induced entanglement (GIE) [${}^{+}$17, MV17] is hypothesizing that not only a massive object can be in a superposition, but that the gravitational field can be used to mediate entanglement.
@@
Note that the two options of either a classical or quantum gravitational correlation (corresponding to gravitational decoherence or GIE, respectively), do not cover the entire field of possibilities for the outcome of the experiment. There is a possibility of quantum gravitational correlation without entanglement. That is, a massive object can be maintained in a coherent superposition, as opposed to gravitational decoherence, but gravity alone cannot create entanglement, as opposed to GIE. Indeed, some preliminary results suggest that even if spacetime is quantum, this does not necessarily implies it has to mediate entanglement [${}^{+}$18, ${}^{+}$19, Ber19, DSW22, SMY23].
@@
The question we focus on in this research is, what are the computational implications of the the last scenario? Assuming quantum gravitational correlations without entanglement, what happens when the system $S_{U}$ executes, while the system $S_{M}$ finished executing and is in the state $\frac{1}{\sqrt{2}}\cdot|{L_{0}}\rangle_{S_{M}}+\frac{1}{\sqrt{2}}\cdot|{L_{1}}\rangle_{S_{M}}$ Our computational model makes the following assumptions. First, if quantum gravitational correlations are possible but cannot generate entanglement, when executing the experiment this should create a joint state that resembles the un-normalized state
@@
We observe that the norm of the state $U_{1}U_{0}|{\psi}\rangle+U_{0}U_{1}|{\psi}\rangle$ fluctuates as a function of all components $U_{0}$, $U_{1}$ and $|{\psi}\rangle$. This means that (similarly to calculations stemming from general relativity in quantum gravity) such transformation will not be linear, due to the need for a state re-normalization. We imagine that in case that the physical universe allows to execute transformations $\left(U_{0},U_{1},|{\psi}\rangle\right)\rightarrow U_{1}U_{0}|{\psi}\rangle+U_{0}U_{1}|{\psi}\rangle$, it is numerically intuitive that decoherence will be inversely proportional to the norm of the state, and that energy could be invested to maintain coherence, which should also be inversely proportional to the norm. The above is the inspiration behind our computational model, which we present next.
@@ 1.2 Our Computational Model - Computable Order Interference
Execution of superposition of unitary orders as a function of interference norm.
@@
On the other hand, for every number of unitaries $m\in\mathbb{N}$ and state $|{\psi}\rangle$, if $U_{1}=U_{2}=\cdots=U_{m}$, one can verify that $m!$ is the norm. The previous work of [Cos22] checks whether (according to the PMF) it is possible to always and unconditionally execute a superposition of unitary orders, for every unitaries $\{U_{i}\}_{i\in[m]}$ and state $|{\psi}\rangle$. In our relaxed model, we have two main differences: We take into consideration computational complexity, and we take into consideration the norm of the OI state, in particular as a part of the computational complexity of an algorithm’s execution.
@@
Our next observation is that when trying to execute OI transformations in standard quantum mechanics, the consideration of the norm is in fact expressed. More formally, we show there is a quantum algorithm $Q$ that given input unitary circuits $\{U_{i}\}_{i\in[m]}$ (all act on the same number of qubits $n$) and an $n$-qubit state $|{\psi}\rangle$, executes in time $\mathsf{poly}\left(\sum_{i\in[m]}|U_{i}|\right)$, where $|U_{i}|$ is the circuit size of $U_{i}$, and outputs the normalized OI state $\textbf{OI}\left(\{U_{i}\}_{i\in[m]},|{\psi}\rangle\right)$ with success probability $\left(\frac{\left\lVert\textbf{OI}\left(\{U_{i}\}_{i\in[m]},|{\psi}\rangle\right)\right\rVert}{m!}\right)^{2}$. We elaborate formally on the above quantum procedure in Section 4.2. There is no known way in quantum mechanics to amplify the probability for obtaining the OI state, other than repeated trials.
@@
Rather than either the bounded interference in quantum mechanics or an unconditional interference, in this work we investigate the computational power of quantum computing with computable interference of unitary orders. In our model, while the probability to succeed in executing order interference is proportional to the norm of the OI state (as per standard quantum mechanics), we allow an investment of computational work (or energy) to amplify the success probability arbitrarily close to $1$. The ability to execute computable order interference is formally captured by an object we define, called an Order Interference (OI) oracle $\mathcal{O}_{\textbf{OI}}$. The full definition (Definition 4.4) of an OI oracle is given in Section 4.1, and for the sake of this Section 1 and Section 2, it will be sufficient to focus on a slightly weaker variant of it (given in Definition 4.2).
@@
and takes time complexity $O\left(\lambda+\sum_{i\in[m]}|U_{i}|\right)$. At the end of execution, with probability
$\frac{\frac{\left\lVert\textbf{OI}\left(\{U_{i}\}_{i\in[m]},|{\psi}\rangle\right)\right\rVert}{m!}}{\frac{\left\lVert\textbf{OI}\left(\{U_{i}\}_{i\in[m]},|{\psi}\rangle\right)\right\rVert}{m!}+\frac{1}{\lambda}}\enspace,$ (3)
@@ 1.3 Results
We focus on two of the most well-studied computational tasks in computer science: The Graph Isomorphism problem ($\mathsf{GI}$, Definition 3.9) and the Gap Closest Vector Problem ($\mathsf{GapCVP}_{g(n)}$, Definition 3.11). Both $\mathsf{GI}$ and $\mathsf{GapCVP}_{g(n)}$ for $g(n):=O(n^{2})$ are believed by experts to be unsolvable in quantum polynomial time. In particular, $\mathsf{GapCVP}_{O(n^{2})}$ is a key problem in the theory of computation and have an immense practical significance: Solving $\mathsf{GapCVP}_{O(n^{2})}$ means solving the Learning With Errors ($\mathsf{LWE}$) problem [Reg09], the computational hardness of which is the central pillar of post-quantum cryptography. We prove the following theorems.
@@
[…]
@@
On a high-level, this work shows a new connection between information processing models in quantum gravity, and some of the fundamental computational problems in computer science. we make the following contributions to computational complexity and quantum information theory.
  • In computational complexity, we show a novel parameterization technique for classical circuits $C:\{0,1\}^{k^{\prime}}\rightarrow\{0,1\}^{k}$, allowing to view invertibility of circuits as a spectrum rather than a binary predicate. We define a new natural and relaxed variant of the known Statistical Difference [SV03] problem ($\mathsf{SD}$, Definition 3.5), called the Sequentially-Invertible Statistical Difference problem ($\mathsf{SISD}$, Definition 4.10). Previous work [GMW91], [GG98] shows how to reduce both $\mathsf{GI}$ and $\mathsf{GapCVP}$ to $\mathsf{SD}$, and we show new reductions to the relaxed problem $\mathsf{SISD}$. These results are in classical computational complexity theory, independent of quantum computation.
  • In quantum information theory, we define a new computational model and complexity measure, capturing a natural analogue of quantum computation with unitary order superposition. We observe a connection between the model and the problem of Statistical Difference $\mathsf{SD}$. As a further technical step, we show a new algorithmic technique in the computational model, which in turn shows how for some parameters, the $\mathsf{SISD}$ problem can be solved by a quantum computer with unitary order superposition (in formal terms, that for some parameters, $\mathsf{SISD}\in\textbf{BQP}^{\text{OI}}$).
@@ 2.1 Computable Order Interference and Quantum State Generation
Before we are trying to use computable OI to solve new problems, it is important to intuitively understand what it enables that a standard quantum computer does not. In the introduction we saw that a standard quantum computer successfully executes an OI transformation with probability $\left(\frac{\left\lVert\textbf{OI}\left(\{U_{i}\}_{i\in[m]},|{\psi}\rangle\right)\right\rVert}{m!}\right)^{2}$, that is, in the order of the scaled norm of the OI state. In the setting of computable OI (modeled by an OI oracle $\mathcal{O}_{\textbf{OI}}$), this probability can be amplified as described in 3. One can verify that to make the success probability 3 at least a constant using computable OI, the complexity of amplification (captured by the parameter $\lambda$) has to be at least the inverse of the scaled norm $\Theta\left(\frac{m!}{\left\lVert\textbf{OI}\left(\{U_{i}\}_{i\in[m]},|{\psi}\rangle\right)\right\rVert}\right)$, depends on the wanted constant [^4] More generally, to get a success probability of $1-\frac{1}{k}$, executing the OI oracle with $\lambda=k\cdot\frac{m!}{\left\lVert\textbf{OI}\left(\{U_{i}\}_{i\in[m]},|{\psi}\rangle\right)\right\rVert}$ suffices.. However, if we are already allowed to execute in the complexity scales of $\frac{m!}{\left\lVert\textbf{OI}\left(\{U_{i}\}_{i\in[m]},|{\psi}\rangle\right)\right\rVert}$, a standard quantum computer can successfully generate the OI state with high probability: Just by repeatedly trying to execute the OI transformation, where as we know, each attempt succeeds with the bounded probability $\left(\frac{\left\lVert\textbf{OI}\left(\{U_{i}\}_{i\in[m]},|{\psi}\rangle\right)\right\rVert}{m!}\right)^{2}$. When an OI transformation fails in either of the cases (computable OI or standard quantum computing) the state is considered as destroyed. So, in both cases, with or without computable OI, the computational complexity of successfully executing a single OI transformation is $\approx\frac{m!}{\left\lVert\textbf{OI}\left(\{U_{i}\}_{i\in[m]},|{\psi}\rangle\right)\right\rVert}$.
@@
What is the complexity of executing $2$ consecutive OI transformations? For two different sets of unitaries $\{U_{0,i}\}_{i\in[m_{0}]}$, $\{U_{1,i}\}_{i\in[m_{1}]}$. Assuming both transformations are successful, this will create a state of different structure than what can be created by a single OI transformation. The gap in complexity between standard quantum computation and computable OI, becomes more apparent when we think of running $\ell$ consecutive OI transformations, for a general $\ell$. Computable OI lets us maintain a probability close to $1$ in that case, using only complexity $\approx\ell^{3}\cdot\frac{m!}{\left\lVert\textbf{OI}\left(\{U_{i}\}_{i\in[m]},|{\psi}\rangle\right)\right\rVert}$ over all $\ell$ attempted transformations, while a standard quantum computer needs all $\ell$ tries to succeed consecutively, which generally has a complexity increasing exponentially with $\ell$. One deduction is thus, that the power of computable OI is by executing OI transformations repeatedly. Assume we are starting from the state $|{0^{n}}\rangle$, and we sequentially execute OI transformations, for different sets of unitaries. The resulting quantum state is
$\sum_{\sigma_{1}\in S_{m_{1}},\cdots,\sigma_{\ell}\in S_{m_{\ell}}}\left(\prod_{i_{\ell}\in[m_{\ell}]}U_{\sigma_{\ell}^{-1}\left(i_{\ell}\right)}\right)\cdots\left(\prod_{i_{1}\in[m_{1}]}U_{\sigma_{1}^{-1}\left(i_{1}\right)}\right)|{\psi}\rangle\enspace,$ (4)
@@
and has a unique structure, corresponding to a uniform superposition over all partial execution orders, where the partial order rule is that for $i,j\in[\ell]$, $i<j$, the execution of unitaries in the $i$-th set $\{U_{i,k}\}_{k\in[m_{i}]}$ always come before the execution of unitaries in the $j$-th set $\{U_{j,k}\}_{k\in[m_{j}]}$.

### v2

@@ Abstract
We show that a quantum computer with the ability to create a superposition of unitary orders is able to solve in polynomial time two fundamental problems in computer science: The Graph Isomorphism Problem ($\mathsf{GI}$) and the Gap Closest Vector Problem ($\mathsf{GapCVP}$), with gap $O\left(n^{2}\right)$. These problems are believed by experts to be hard to solve for a regular quantum computer. Interestingly, our model does not seem overpowered, and we found no obvious way to solve entire complexity classes that are considered hard in computer science, like the classes NP and SZK. As part of our work we develop a new parameterization technique for the invertibility of classical circuits, which raises new questions in the analysis of probability distributions and induces a new parameter for hardness inside the complexity class SZK. These new techniques are independent of any non-classical computation.
@@ 1 Introduction
Formulating an experimentally-verified theory of physics that generalizes both QM and GR, is one of the grand challenges of physics. Such a theory, usually referred to as a theory of Quantum Gravity, will model the interaction between quantum effects and gravitational force; an interaction which is currently ignored in most analyses. The two most popular candidates as theories of quantum gravity are String Theory and Loop Quantum Gravity, however, numerous other proposals exist [dBDE+22].
@@
The goal of this paper is to suggest and define a computational model for a quantum computer with the added ability of generating a superposition of spacetime geometries, and to analyze its power. The rest of the Introduction is as follows. In Section 1.1 we provide a brief overview of the physics literature on a superposition of spacetime geometries, and then explain how it serves as inspiration for our computational model. We also provide the physical assumptions our model makes. Section 1.1 of the introduction deals with physics, it is self contained and can be skipped for the readers that are interested only in the complexity theoretical aspects of this work. In Section 1.2 we present our computational model and its connection to both standard quantum computation and previous generalized models of quantum information processing. In Section 1.3 we present our complexity theoretical results for the model.
@@ 1.1 Gravity, Superposition and Entanglement
Before we explain what a quantum spacetime may be able to do, we refer to one ability of a classical spacetime. GR postulates the effects of gravitational time dilation, where the structure of spacetime determines the pace at which time flows. More precisely, gravitational time dilation is the effect of time moving slower in a region of space with greater gravitational potential: For some location $s\in\mathbb{R}^{3}$ in space, the time in $s$ slows down proportionally to how close $s$ is to a massive object, and the mass of that object. According to this logic, a quantum spacetime may enable a superposition of spacetime curvatures, which in turn induces a superposition of time dilation effects, on the same set of events.
@@
  • The last detail is the only one that uses effects from GR, specifically, gravitational time dilation. The system $S_{M}$ is located with respect to the second system $S_{U}$ such that, even though the systems are spatially disjoint in coordinates, if $M$ is at location $L_{0}$, which is closer to clock $C_{1}$, then $C_{1}$ ticks slower and $C_{0}$ concludes first, and if $M$ is at location $L_{1}$, which is closer to $C_{0}$, then $C_{0}$ ticks slower and $C_{1}$ concludes first.
Imagine we (1) execute $S_{M}$ and generate a uniform superposition of the two possible locations for $M$, followed by (2) executing $S_{U}$, the clocked execution of the unitaries $U_{0}$, $U_{1}$. We can analyze what happens in each of the branches of the superposition, separately, assuming as if spacetime is classical and there is no superposition of the locations of $M$. If $M$ is at $L_{0}$, the state in $R$ at the end of the execution is $U_{1}U_{0}|{\psi}\rangle$, and if $M$ is at $L_{1}$ then the state in $R$ at the end of execution is $U_{0}U_{1}|{\psi}\rangle$. The question becomes one in quantum gravity, when we ask what happens not in each branch of the superposition separately, but as a full joint system, in superposition. Experimental physics does not know the answer to this question.
The above question may translate to more basic questions about gravity, superposition and entanglement. Specifically, the first question is whether a massive object can be maintained in a quantum superposition, and the second question is whether entanglement can be created through gravity alone. To elaborate, on one side, the gravitational decoherence hypothesis [Dió89, Pen96, Dio14, Dio87] says that a quantum superposition of a massive object will fundamentally decohere, turning into a probabilistic mixture of classical states. On the other end, gravitationally-induced entanglement (GIE) [BMM+17, MV17] is hypothesizing that not only a massive object can be in a superposition, but that the gravitational field can be used to mediate entanglement.
@@
Note that the two options of either a classical or quantum gravitational correlation (corresponding to gravitational decoherence or GIE, respectively), do not cover the entire field of possibilities for the outcome of the experiment. There is a possibility of quantum gravitational correlation without entanglement (quantum correlations are known to be possible without entanglement, e.g., quantum discord [OZ01]). That is, a massive object can be maintained in a coherent superposition, as opposed to gravitational decoherence, but gravity alone cannot create entanglement, as opposed to GIE. Indeed, some preliminary results suggest that even if spacetime is quantum, this does not necessarily implies it has to mediate full entanglement [BWG+18, BWG+19, Ber19, DSW22, SMY23].
@@
The question we focus on in this research is, what are the computational implications of the the last scenario? Assuming quantum gravitational correlations without (full) entanglement, what happens when the system $S_{U}$ executes, while the system $S_{M}$ finished executing and is in the state $\frac{1}{\sqrt{2}}\cdot|{L_{0}}\rangle_{S_{M}}+\frac{1}{\sqrt{2}}\cdot|{L_{1}}\rangle_{S_{M}}$ Our computational model makes the following assumptions. First, if quantum gravitational correlations are possible but do not force entanglement, when executing the experiment, this should create a joint state that resembles the un-normalized state
@@
We observe that the norm of the state $U_{1}U_{0}|{\psi}\rangle+U_{0}U_{1}|{\psi}\rangle$ fluctuates as a function of all components $U_{0}$, $U_{1}$ and $|{\psi}\rangle$. This means that (similarly to calculations stemming from general relativity in quantum gravity) such transformation will not be linear, due to the need for a state re-normalization. Next, we imagine that in case that the physical universe allows to execute transformations $\left(U_{0},U_{1},|{\psi}\rangle\right)\rightarrow U_{1}U_{0}|{\psi}\rangle+U_{0}U_{1}|{\psi}\rangle$, it is numerically intuitive that decoherence will be inversely proportional to the norm of the state, and that energy could be invested to maintain coherence, which should also be inversely proportional to the norm. Finally, while we assume the ability to invest energy to maintain coherence of a superposition of spacetimes, we do not assume that matter-wave interference between different spacetimes to be necessarily identical to quantum interference within a single spacetime [^2] Some theoretical results show evidence that standard interference dynamics is not necessarily the case when considering a superposition of spacetimes [FMZ21, FAZM22, FAZM23].. Later when we define our computational model, these assumptions will be more formally expressed.
@@ 1.2 Our Computational Model - Computable Order Interference
Execution of superposition of unitary orders as a function of interference norm and phase alignment.
@@
On the other hand, for every number of unitaries $m\in\mathbb{N}$ and state $|{\psi}\rangle$, if $U_{1}=U_{2}=\cdots=U_{m}$, one can verify that $m!$ is the norm.
The previous work of [Cos22] checks whether (according to the PMF) it is possible to always and unconditionally execute a superposition of unitary orders, for every unitaries $\{U_{i}\}_{i\in[m]}$ and state $|{\psi}\rangle$. In our relaxed model, we have the following main differences. We take into consideration computational complexity, and specifically, the computational complexity of generating the OI state is determined by two properties of the interference [^5] These complexity measures are tied to the physical assumptions given at the end of Section 1.1.: (1) The norm of the OI state, and (2) The amount of phase alignment between execution orders (or spacetimes). More precisely, for every classical state $|{x}\rangle$ (for $x\in\{0,1\}^{n}$), we consider the phase and magnitude of $|{x}\rangle$ in the resulting state $\left(\prod_{i\in[m]}U_{\sigma^{-1}\left(i\right)}\right)\cdot|{\psi}\rangle$ of every execution order $\sigma\in S_{m}$. The amount of phase alignment is how similar are the phases of the same $|{x}\rangle$, between the results of differing execution orders, over all $x\in\{0,1\}^{n}$.
@@
Our next observation is that when trying to execute OI transformations in standard quantum mechanics, the consideration of the norm is in fact expressed (a limited phase alignment implies reduced norm, thus phase alignment possibly also plays a role). More formally, we show there is a quantum algorithm $Q$ that given input unitary circuits $\{U_{i}\}_{i\in[m]}$ (all act on the same number of qubits $n$) and an $n$-qubit state $|{\psi}\rangle$, executes in time $\mathsf{poly}\left(\sum_{i\in[m]}|U_{i}|\right)$, where $|U_{i}|$ is the circuit size of $U_{i}$, and outputs the normalized OI state $\textbf{OI}\left(\{U_{i}\}_{i\in[m]},|{\psi}\rangle\right)$ with success probability $\left(\frac{\left\lVert\textbf{OI}\left(\{U_{i}\}_{i\in[m]},|{\psi}\rangle\right)\right\rVert}{m!}\right)^{2}$. We elaborate formally on the above quantum procedure in Section 4.2. There is no known way in quantum mechanics (and in particular, in the standard quantum circuit model) to generally and efficiently amplify the success probability of executing an OI transformation.
@@
Rather than either the bounded interference in quantum mechanics or an unconditional interference, in this work we investigate the computational power of quantum computing with computable interference of unitary orders. In our model, while the probability to succeed in executing order interference is proportional to the norm of the OI state (as per standard quantum mechanics), we allow an investment of computational work (or energy) to amplify the success probability arbitrarily close to $1$ (up to phase misalignment). The ability to execute computable order interference is formally captured by an object we define, called an Order Interference (OI) oracle $\mathcal{O}_{\textbf{OI}}$. The full definition (Definition 4.4) of an OI oracle is given in Section 4.1, and for the sake of this Section 1 and Section 2, it will be sufficient to focus on a relaxed variant of it (given in a Corollary 4.1 of the definition).
@@
and takes time complexity $O\left(\lambda+\sum_{i\in[m]}|U_{i}|\right)$. For $x\in\{0,1\}^{n}$, $\sigma\in S_{m}$ we write
$\prod_{i\in[m]}U_{\sigma^{-1}\left(i\right)}\cdot|{\psi}\rangle:=\sum_{x\in\{0,1\}^{n}}\alpha_{x,\sigma}|{x}\rangle\enspace.$
At the end of execution, with probability
$\left(\frac{\sum_{x\in\{0,1\}^{n}}|\sum_{\sigma\in S_{m}}\alpha_{x,\sigma}|}{\sum_{x\in\{0,1\}^{n}}\sum_{\sigma\in S_{m}}|\alpha_{x,\sigma}|}\right)\frac{\frac{\left\lVert\textbf{OI}\left(\{U_{i}\}_{i\in[m]},|{\psi}\rangle\right)\right\rVert}{m!}}{\frac{\left\lVert\textbf{OI}\left(\{U_{i}\}_{i\in[m]},|{\psi}\rangle\right)\right\rVert}{m!}+\frac{1}{\lambda}}\enspace,$ (3)
@@ 1.3 Results
We focus on two very well-studied computational tasks in computer science: The Graph Isomorphism problem ($\mathsf{GI}$, Definition 3.9) and the Gap Closest Vector Problem ($\mathsf{GapCVP}_{g(n)}$, Definition 3.11). Both $\mathsf{GI}$ and $\mathsf{GapCVP}_{g(n)}$ for $g(n):=O(n^{2})$ are believed by experts to be unsolvable in quantum polynomial time. In particular, $\mathsf{GapCVP}_{O(n^{2})}$ is a key problem in the theory of computation and have an immense practical significance: Solving $\mathsf{GapCVP}_{O(n^{2})}$ means solving the Learning With Errors ($\mathsf{LWE}$) problem [Reg09], the computational hardness of which is the central pillar of post-quantum cryptography. We prove the following theorems.
@@
There exists a positive absolute constant $c\in\mathbb{R}_{>0}$ such that $\mathsf{GapCVP}_{c\cdot n^{2}}\in\textbf{BQP}^{\text{OI}}$.
@@
As a high-level summary, this work shows a new connection between information processing models in quantum gravity, and some of the fundamental computational problems in computer science. More specifically, we make the following contributions to computational complexity and quantum information theory (both are explained in length in Section 2).
  • In computational complexity, we show a novel parameterization technique for classical circuits $C:\{0,1\}^{k^{\prime}}\rightarrow\{0,1\}^{k}$, allowing to view invertibility of circuits as a spectrum rather than a binary predicate. We define a new natural and relaxed variant of the known Statistical Difference [SV03] problem ($\mathsf{SD}$, Definition 3.5), called the Sequentially-Invertible Statistical Difference problem ($\mathsf{SISD}$, Definition 4.12). Previous work [GMW91], [GG98] shows how to reduce both $\mathsf{GI}$ and $\mathsf{GapCVP}$ (respectively) to $\mathsf{SD}$, and as a technical contribution we show new reductions to the relaxed problem $\mathsf{SISD}$. These results are in classical computational complexity theory, independent of any non-classical physics or computation.
  • In quantum information theory, we define a new computational model and complexity measure, capturing a simple analogue of quantum computation with unitary order superposition. We observe a connection between the model and the problem of Statistical Difference $\mathsf{SD}$. As a further technical step, we show a new algorithmic technique in the computational model, which in turn implies that for some parameters, the $\mathsf{SISD}$ problem can be solved by a quantum computer with unitary order superposition (in formal terms, that for some parameters, $\mathsf{SISD}\in\textbf{BQP}^{\text{OI}}$).
@@ 2.1 Computable Order Interference and Quantum State Generation
Before we are trying to use computable OI to solve new problems, it is important to intuitively understand what it enables that a standard quantum computer does not. In the introduction we saw that a standard quantum computer successfully executes an OI transformation with probability $\left(\frac{\left\lVert\textbf{OI}\left(\{U_{i}\}_{i\in[m]},|{\psi}\rangle\right)\right\rVert}{m!}\right)^{2}$, that is, in the order of the scaled norm of the OI state. In the setting of computable OI (modeled by an OI oracle $\mathcal{O}_{\textbf{OI}}$), this probability can be amplified as described in 3, depending on the amount of phase alignment (i.e., how much the non-amplifiable, first component in 3, is close to $1$). One can verify that to make the success probability 3 at least a constant using computable OI, the complexity of amplification (captured by the parameter $\lambda$) has to be at least the inverse of the scaled norm $\Theta\left(\frac{m!}{\left\lVert\textbf{OI}\left(\{U_{i}\}_{i\in[m]},|{\psi}\rangle\right)\right\rVert}\right)$, depends on the wanted constant [^6] More generally, assuming perfect phase alignment, to get a success probability of $1-\frac{1}{k}$, executing the OI oracle with $\lambda=k\cdot\frac{m!}{\left\lVert\textbf{OI}\left(\{U_{i}\}_{i\in[m]},|{\psi}\rangle\right)\right\rVert}$ suffices.. However, if we are already allowed to execute in the complexity scales of $\frac{m!}{\left\lVert\textbf{OI}\left(\{U_{i}\}_{i\in[m]},|{\psi}\rangle\right)\right\rVert}$, a standard quantum computer can successfully generate the OI state with high probability: Just by repeatedly trying to execute the OI transformation, where as we know, each attempt succeeds with the bounded probability $\left(\frac{\left\lVert\textbf{OI}\left(\{U_{i}\}_{i\in[m]},|{\psi}\rangle\right)\right\rVert}{m!}\right)^{2}$. When an OI transformation fails in either of the cases (computable OI or standard quantum computing) the state is considered as destroyed. So, in both cases, the computational complexity of successfully executing a single OI transformation is $\approx\frac{m!}{\left\lVert\textbf{OI}\left(\{U_{i}\}_{i\in[m]},|{\psi}\rangle\right)\right\rVert}$.
@@
What is the complexity of executing $2$ consecutive OI transformations? For two different sets of unitaries $\{U_{0,i}\}_{i\in[m_{0}]}$, $\{U_{1,i}\}_{i\in[m_{1}]}$. Assuming both transformations are successful, this will create a state of different structure than what can be created by a single OI transformation. The gap in complexity between standard quantum computation and computable OI, becomes more apparent when we think of running $\ell$ consecutive OI transformations, for a general $\ell$ (assuming perfect phase alignment in both cases). Computable OI lets us maintain a probability close to $1$ in that case, using only complexity $\approx\ell^{3}\cdot\frac{m!}{\left\lVert\textbf{OI}\left(\{U_{i}\}_{i\in[m]},|{\psi}\rangle\right)\right\rVert}$ over all $\ell$ attempted transformations, while a standard quantum computer needs all $\ell$ tries to succeed consecutively, which generally has a complexity increasing exponentially with $\ell$. One deduction is thus, that the power of computable OI is by executing OI transformations sequentially. Assume we are starting from the state $|{0^{n}}\rangle$, and we sequentially execute $\ell$ OI transformations, for $\ell$ different sets of unitaries $\{U_{1,i}\}_{i\in[m_{1}]}$, ,$\cdots,\{U_{\ell,i}\}_{i\in[m_{\ell}]}$. The resulting quantum state is
$\sum_{\sigma_{1}\in S_{m_{1}},\cdots,\sigma_{\ell}\in S_{m_{\ell}}}\left(\prod_{i_{\ell}\in[m_{\ell}]}U_{\ell,\sigma_{\ell}^{-1}\left(i_{\ell}\right)}\right)\cdots\left(\prod_{i_{1}\in[m_{1}]}U_{1,\sigma_{1}^{-1}\left(i_{1}\right)}\right)|{\psi}\rangle\enspace,$ (4)
@@
and has a unique structure, corresponding to a uniform superposition over all partial execution orders, where the partial order rule is that for $i,j\in[\ell]$, $i<j$, the execution of unitaries in the $i$-th set $\{U_{i,k}\}_{k\in[m_{i}]}$ always comes before the execution of unitaries in the $j$-th set $\{U_{j,k}\}_{k\in[m_{j}]}$.